car A is going 50 mph and car B is going 60 mph and car A is 15 miles ahead how long will it take Car B to catch up
Answer by Edwin McCravy(20086)

The approach rate, or "catch-up rate" is the difference in speeds,
and 60-50 is 10 mph.  To catch up 15 miles at a rate of 10 mph will
take t = d/r = 15/10 = 1.5 hours, an hour and a half.

Edwin